OpenStudy (anonymous):

what is 12 over 9 minus 6 I fraction form?

OpenStudy (anonymous):

To add or subtract fractions the denominator of both fractions must be the same. You can make a whole number into a fraction by multiplying the numerator and denominator by the same number. ( 6 = (6/1) = (6*5/1*5) = (30/5) ) After you have found a common denominator you can simply add the numerators from each fraction together and keep the denominator the same.